An orthorhombic crystal has the following
parameters :
a = 8.2 x 10-10m
b = 9.4 x 10-10m
c = 7.5 x 10-10m
Find the distance between (123) planes.
1
Expert's answer
2019-05-31T00:40:06-0400

Solution.

"\frac{h^2+k^2}{a^2} + \frac{l^2}{c^2} = \frac{1}{d^2}"

h = 1

k = 2

l = 3


"\frac{(h^2+k^2) \times c^2 + (l \times a)^2}{(a \times c)^2} = \frac{1}{d^2}"

"d = 2.07 \times 10^{-10}"

Answer:

"d = 2.07 \times 10^{-10}"